OpenStudy (anonymous):

20 is 4% of what? a.50 b.200 c.500 d.800

OpenStudy (anonymous):

(4/100)*x=20 solve for x

OpenStudy (anonymous):

d.800?

OpenStudy (anonymous):

no x=20*100/4=500

OpenStudy (anonymous):

So, C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thaaanks.

OpenStudy (candy13106):

ccccccccccccc

OpenStudy (candy13106):